If a is equal to 1.2 ± 0.1cm, b=2.0 ± 0.2cm, c=1.05±0.05cm, calculate the following with the errors and standard errors





(i) ab/C³





(ii) ABC





(iii) C²a³/b²

1
Expert's answer
2022-03-06T15:17:41-0500

1)

"\frac{ab}{c^3}=2.07±0.04,"

2)

"abc=2.52\u00b10.35,"

3)

"\frac{c^2a^3}{b^2}=0.47±0.12."